OpenStudy (anonymous):

Solve Radical x + 14 - 12 = x

OpenStudy (anonymous):

check your equation

OpenStudy (anonymous):

\[\sqrt{x + 14} - 12 = x\]

OpenStudy (anonymous):

just take 12 right side and square both the sides

OpenStudy (anonymous):

you will get a quadratic equation

OpenStudy (anonymous):

calculate its roots

OpenStudy (anonymous):

is that ok for u

OpenStudy (australopithecus):

\[(x+14)^{\frac{1}{2}} = x + 12\] \[(x+14)^{\frac{1}{2}*2} = (x + 12)^2\] \[(x+14)^1 = (x + 12)^2\] \[x+14 = (x + 12)^2\]

OpenStudy (anonymous):

My Choices Are X = -10 X - -13 X =-10 & -13 And None Of The Above . . . So Im Assuming Its None Of The Above

OpenStudy (australopithecus):

dont assume

OpenStudy (australopithecus):

do you not follow my algebra

OpenStudy (australopithecus):

Just look at it and try to understand

OpenStudy (australopithecus):

do you not know how to expand \[x + 14 = (x+12)(x+12)\]

OpenStudy (australopithecus):

(a+b)(a+b) = a^2 + ab + ab + bb